# Large language model validity via enhanced conformal prediction methods

### v1

@@ Abstract
We develop new conformal inference methods for obtaining validity guarantees on the output of large language models (LLMs). Prior work in conformal language modeling identifies a subset of the text that satisfies a high-probability guarantee of correctness. These methods work by filtering claims from the LLM’s original response if a scoring function evaluated on the claim fails to exceed a threshold calibrated via split conformal prediction. Existing methods in this area suffer from two deficiencies. First, the guarantee stated is not conditionally valid. The trustworthiness of the filtering step may vary based on the topic of the response. Second, because the scoring function is imperfect, the filtering step can remove many valuable and accurate claims. We address both of these challenges via two new conformal methods. First, we generalize the conditional conformal procedure of Gibbs et al. (2023) in order to adaptively issue weaker guarantees when they are required to preserve the utility of the output. Second, we show how to systematically improve the quality of the scoring function via a novel algorithm for differentiating through the conditional conformal procedure. We demonstrate the efficacy of our approach on both synthetic and real-world datasets.
@@ 1 Introduction
Since an LLM output is not always trustworthy, a growing body of work aims to quantify uncertainty regarding a given output’s validity. While there are many approaches to this problem [26, 4, 9, 16], this paper considers a particularly popular method for black-box uncertainty quantification: conformal inference [28, 2, 3]. Conformal inference provides a generic procedure for transforming the predictions of any modeling procedure into valid prediction sets that are guaranteed to contain the true outcome with high probability. Several recent papers have applied conformal inference to define a set of LLM responses that contains at least one factual response with high probability [3, 15, 23, 30]. But while generating a candidate set of outputs may be a reasonable strategy in some question-answering problems, it is not a generalizable approach for the diverse and unstructured tasks faced in real-world deployment.
@@
Figure 1: The left panel displays the output of GPT-3.5-Turbo for the prompt “How often is a shingles vaccine required?” The first filtered output (center) is calibrated using the frequency score described in Section D.1. The second filtered output (right) is calibrated using a score obtained via conditional boosting. Both filtered outputs are guaranteed to include no false claims with the stated probability. The middle panel displays the output of the marginally valid conformal factuality method of Mohri and Hashimoto [19] at a fixed level of 90%. The rightmost panel displays the output of our method at a level of 63%, which is chosen adaptively to ensure that at least 70% of the claims are retained. We calibrate our method to this adaptive level using a generalization of conditional conformal prediction [10].
@@ 1.1 Summary of contributions
Our first method, which we call conditional boosting, allows for the automated discovery of superior claim scoring functions via differentiation through the conditional conformal algorithm of Gibbs et al. [10]. Automated conformal score improvement was introduced by Stutz et al. [25]; their paper shows how to minimize conformal prediction set size in a classification setting by differentiating through the split conformal algorithm. As we show, however, in Section 4, optimizing the score function subject only to a marginal coverage constraint can lead to poor conditional properties.
Differentiating through the conditional conformal algorithm presents new challenges that are, to the best of our knowledge, unaddressed by prior work. Our key technical contributions are a proof that the cutoff output by the conditional conformal method remains locally linear with respect to the conformity score under mild assumptions as well as a computationally efficient estimator of this derivative. By running gradient descent with this derivative, we discover new scores that enable greater claim retention.
The right panel of Figure 2 demonstrates the efficacy of our method. Here, we use boosting to learn an optimal linear combination of four candidate scoring functions. We compare the learned, boosted scores (orange) against a baseline method (blue) that uses the “frequency” scoring method developed by Mohri and Hashimoto [19]. As the figure shows, the boosted score allows for higher claim retention across all datasets (mean claim retention of 39% vs. 24% for the boosted vs. unboosted scores).
Our second method, which we call level-adaptive conformal prediction, allows the validity of the conformal output to depend on characteristics of the queried prompt. In our LLM experiments, we adapt the level, i.e., the claimed probability of correctness, to each prompt in order to ensure that issued outputs retain at least $70\%$ of the original set of sub-claims. For example, in Figure 1, we prompt GPT-3.5-Turbo to output a response to a question from the MedicationQA dataset [6]. Outputting a filtered response that achieves the stated factuality criterion with probability $90\%$ requires near-complete censorship, but by relaxing the level to $63\%$ using our method, we can preserve almost the entire response.
@@
Figure 2 displays the advantages of our approach to this problem. First, the left panel of Figure 2 verifies that the level-adaptive probabilities we report are empirically well-calibrated. Second, the right panel of Figure 2 quantitatively demonstrates the improved claim retention of our method. For each dataset included in the MedLFQA benchmark [13], we show that level-adaptive conformal prediction retains at least 70% of the original output’s claims in most examples. By combining our two methods, we retain most claims and output non-trivial guarantees of response factuality; the left panel shows that the issued probabilities vary between $50$ and $85\%$. By contrast, while the fixed level method guarantees a $90\%$ probability of correctness, the method retains very little of the original LLM output.
@@
More concretely, (1) establishes that the issued probabilities are well-calibrated in the following sense: among similar prompts, the outputs that we claim to be factually correct with probability, say, between 70 and 80% will be actually factual between 70 and 80% of the time. This guarantee holds in expectation over the dataset used to run our procedure. In this paper, we also allow for many notions of “factuality,” e.g., in Figure 1, we guarantee that filtered outputs have no unsubstantiated claims, but in other examples, we relax our guarantee so that no more than $k$ false claims are issued.
@@
We release a filtered version of the MedLFQA benchmark that removes some non-health-related prompts, the generated and parsed text used to run our experiments, as well as the notebooks used to produce the figures in this paper at github.com/jjcherian/conformal-safety. We also update our Python package for conditional conformal inference to support level-adaptive conformal prediction. The new version of this package is available to download at github.com/jjcherian/conditional-conformal and can be installed from PyPI.
@@ 2.1 Conformal prediction with conditional guarantees
To make this guarantee concrete, we might define :=$\mathcal{F}:=\{(\mathbbm{1}\{X\in G\})^{\top}\beta\mid\beta\in\mathbb{R}^{|\mathcal{G}|}\}$ for some set of subgroups $\mathcal{G}$. Then, (3) is exactly equivalent to group-conditional coverage, i.e., $\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})\mid X_{n+1}\in G)=1-\alpha$ for all $G\in\mathcal{G}$.
To understand their construction, let $S(X,Y)\in\mathbb{R}$ denote a conformity score that measures how well $Y$ “conforms” to an estimate of its value. A typical choice in the regression setting is $S(X,Y)=|Y-\hat{\mu}(X)|$ for some fixed regression function $\hat{\mu}(\cdot)$; in the classification setting, we might choose $S(X,Y)=1/\hat{\pi}(Y\mid X)$ for some estimated conditional probabilities $\hat{\pi}$. Then, Gibbs et al. [10] estimate an upper bound for these scores by fitting an augmented quantile regression in which the unknown test score, $S(X_{n+1},Y_{n+1})$ is replaced by an imputed value $S$. Formally, let $\ell_{\alpha}(r):=(1-\alpha)[r]_{+}+\alpha[r]_{-}$ denote the pinball loss. Then, they define
@@
Since $\hat{C}(X_{n+1})$ can be conservative, Gibbs et al. [10] also define a smaller randomized prediction set, $\hat{C}_{\text{rand.}}(X_{n+1})$ that achieves exact coverage. We will typically prefer to work with this set, but defer a detailed definition of its construction to the Appendix. As the following theorem shows, this randomized set achieves the guarantee stated in (3).
@@ 2.2 Conformal factuality
As discussed in the introduction, prediction sets are not a practical object for LLM outputs. Thus, to adapt conformal inference to this setting, Quach et al. [23] and Mohri and Hashimoto [19] develop filtering methods that remove false claims from an LLM’s output. We will focus specifically here on the work of Mohri and Hashimoto [19] since this is most similar to the new methods that we will propose in this article. Recall that $\mathbf{C}_{i}=\{C_{ij}\}_{j=1}^{k_{i}}$ denotes the claims made in an LLM’s response and $\mathbf{W}_{i}=\{W_{ij}\}_{j=1}^{k_{i}}$ denotes binary variables for which $W_{ij}=1$ indicates that the claim is true.
@@
where $\hat{\tau}$ is the $\frac{\lceil(1-\alpha)(n+1)\rceil}{n+1}$-quantile of the conformity scores
@@
Theorem 1 of [19] shows that if $\{({P}_{i},R_{i},\mathbf{C}_{i},\mathbf{W}_{i})\}_{i=1}^{n+1}$ are exchangeable, this method satisfies (5).
@@ 3.1 Generalization to alternative targets
Our first contribution in this paper will be to generalize the conditional framework of Gibbs et al. [10] to generic alignment tasks. More formally, suppose we are given a loss function $L(\hat{F}(\mathbf{C}_{i}),\mathbf{W}_{i})$ that measures the quality of the filtered output $\hat{F}(\mathbf{C}_{i})$ relative to the ground truth $\mathbf{W}_{i}$. Our goal will be to ensure that $\mathbb{P}(L(\hat{F}(\mathbf{C}_{n+1}),\mathbf{W}_{n+1})\leq\lambda)\geq 1-\alpha$, for some user-specified tolerance $\lambda>0$. For example, in the prior section, $L(\cdot,\cdot)$ was the binary indicator that $\hat{F}(\mathbf{C}_{n+1})$ contains a false claim. More generally, $L(\cdot,\cdot)$ could measure the presence of toxic or discriminatory content in the response.
To incorporate general losses into the conditional conformal framework, we require two assumptions. First, we assume that the method is always permitted to abstain from issuing a response and thus $L(\emptyset,\cdot)=0$. Second, we assume the loss is monotone. Namely, for any sets of claims $\hat{F}_{1}(\mathbf{C}_{i})\subseteq\hat{F}_{2}(\mathbf{C}_{i})$, it must be the case that $L(\hat{F}_{1}(\mathbf{C}_{i}),W_{i})\leq L(\hat{F}_{2}(\mathbf{C}_{i}),W_{i})$. This requirement also means that we only incur loss for outputting incorrect information and do not receive any reward for making a larger number of claims. In the sequel, we will discuss methods for optimizing the filtering procedure to satisfy other auxiliary objectives such as outputting a richer response.
With these assumptions, we extend the calibration procedure of Gibbs et al. [10] as follows. Let $X_{n+1}=X_{n+1}(P_{i},R_{i})$ denote a set of features computed using the prompt and response. Matching (6), we define $p(P_{i},C_{ij})$ to be a score measuring the quality of claim $C_{ij}$. We define the filtered set of claims by $\hat{F}(\mathbf{C}_{i})=F(\mathbf{C}_{i};\hat{\tau}):=\{C_{ij}:p(P_{i},C_{ij})\geq\hat{\tau}\}$, and the conformity score via
$S(\mathbf{C}_{i},\mathbf{W}_{i}):=\inf\{\tau\mid L(F(\mathbf{C}_{i};\tau),\mathbf{W}_{i})\leq\lambda\},$
i.e., the minimum filtering threshold that controls the loss. Then, we set
@@
Let =$\mathcal{F}=\{\Phi(X)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$ denote any finite dimensional linear class. Assume that $\{\mathbf{D}_{i}\}_{i=1}^{n+1}$ are exchangeable, that solutions to (7) and its dual are computed symmetrically in the input data, $L(\cdot,\cdot)$ is monotone, and $L(\emptyset,\cdot)=0$. Then, for all $f\in\mathcal{F}$,
@@
As above, we can make this guarantee concrete by choosing $\mathcal{F}$ to be a linear combination of group indicators, i.e., =$\mathcal{F}=\{(\mathbbm{1}\{X\in G\})^{\top}\beta\mid\beta\in\mathbb{R}^{|\mathcal{G}|}\}$. For example, in our Wikipedia biographies experiment, groups might be defined using the subject’s country of origin. This choice of $\mathcal{F}$ yields
@@ 3.2 Level-adaptive conformal prediction
In addition to adapting the filtering cutoff to $X_{n+1}$, it may also be beneficial to adjust the desired error probability $\alpha$ [^3] Although our methods’ implementations differ, we remark here that the level-adaptive guarantee is inspired by a similar result in the unpublished work of Zhang and Candès [31]. More formally, suppose now that we are given some function, $\alpha(\cdot)$ and our goal is to ensure that $L(\hat{F}(\mathbf{C}_{n+1}),\mathbf{W}_{n+1})\leq\lambda$ with probability $1-\alpha(X_{n+1})$. Recall that in the previous section, we use the pinball loss, $\ell_{\alpha}(\cdot)$ to learn the $1-\alpha$ quantile of $S(\mathbf{C}_{i},\mathbf{W}_{i})$. To control the error rate adaptively, here we will use a data-dependent loss for the $i$-th point, $\ell_{\alpha(X_{i})}(\cdot)$. Then, similar to previous section, we define
@@
and apply filtering at the cutoff $\hat{\tau}_{\text{l.a.}}(X_{n+1}):=\sup\{S:S\leq g_{S}(X_{n+1})\}$. As in the previous section, it is more convenient to work with a slightly smaller randomized cutoff, $\hat{\tau}_{\text{l.a., rand.}}$. The following theorem shows that our previous guarantee for fixed $\alpha$ extends to this new setting.
@@
While this choice of function class elicits an interpretable guarantee, it can be quite large in practice. This can both slow down the computation of the conformal cutoff and reduce the proportion of retained claims. The practitioner has some flexibility in choosing the function class to trade-off between quality of guarantee and computational efficiency / claim retention. For example, this class might be chosen to exactly satisfy other popular (and more efficient) approximations to multicalibration such as low-degree multicalibration [11]. In Section B.2, we briefly explore the consequences of choosing an insufficiently complex function class in a synthetic regression example, but remark that a much more substantial exposition of these trade-offs can be found in Gibbs et al. [10].

### v2

@@ Abstract
We develop new conformal inference methods for obtaining validity guarantees on the output of large language models (LLMs). Prior work in conformal language modeling identifies a subset of the text that satisfies a high-probability guarantee of correctness. These methods work by filtering claims from the LLM’s original response if a scoring function evaluated on the claim fails to exceed a threshold calibrated via split conformal prediction. Existing methods in this area suffer from two deficiencies. First, the guarantee stated is not conditionally valid. The trustworthiness of the filtering step may vary based on the topic of the response. Second, because the scoring function is imperfect, the filtering step can remove many valuable and accurate claims. We address both of these challenges via two new conformal methods. First, we generalize the conditional conformal procedure of Gibbs et al. (2023) in order to adaptively issue weaker guarantees when they are required to preserve the utility of the output. Second, we show how to systematically improve the quality of the scoring function via a novel algorithm for differentiating through the conditional conformal procedure. We demonstrate the efficacy of our approach on biography and medical question-answering datasets.
@@ 1 Introduction
Since an LLM output is not always trustworthy, a growing body of work aims to quantify uncertainty regarding a given output’s validity. While there are many approaches to this problem [28, 4, 9, 18], this paper considers a particularly popular method for black-box uncertainty quantification: conformal inference [30, 2, 3]. Conformal inference provides a generic methodology for transforming the predictions of any modeling procedure into valid prediction sets that are guaranteed to contain the true outcome with high probability. Several recent papers have applied conformal inference to define a set of LLM responses that contains at least one factual response with high probability [3, 17, 25, 32]. But while generating a candidate set of outputs may be a reasonable strategy in some question-answering problems, it is not a generalizable approach for the diverse and unstructured tasks faced in real-world deployment.
@@
Figure 1: The left panel displays the output of GPT-3.5-Turbo for the prompt “How often is a shingles vaccine required?” The first filtered output (center) is calibrated using the frequency score (see Section E.1) and the marginally valid conformal factuality method of Mohri and Hashimoto [21] at a fixed level of 90%. The second filtered output (right) is calibrated using a score obtained via our conditional boosting procedure (Section 3.3) at a level of 63%, which is chosen and calibrated using our adaptive method (Section 3.2) to approximately ensure that at least 70% of the claims are retained. Both filtered outputs are guaranteed to include no false claims with the stated probability.
@@ 1.1 Summary of contributions
Our first method, which we call conditional boosting, allows for the automated discovery of superior claim scoring functions via differentiation through the conditional conformal algorithm of Gibbs et al. [10]. Automated conformal score improvement was introduced by Stutz et al. [27]; their paper shows how to minimize conformal prediction set size in a classification setting by differentiating through the marginally valid split conformal algorithm. As we show, however, in Section 4, optimizing the score function subject only to a marginal coverage constraint can lead to poor conditional properties.
Optimizing through the conditional conformal algorithm is not straightforward. Our key technical contributions are a proof that (under mild assumptions) the cutoff output by the conditional conformal method is differentiable and a computationally efficient method for computing this derivative. By running gradient descent using this algorithm we discover new scores that enable greater claim retention.
The right panel of Figure 2 demonstrates the efficacy of our method. Here, we use boosting to learn an optimal linear combination of four candidate scoring functions. We compare the learned, boosted scores (orange) against a baseline method (blue) that uses the “frequency” scoring method developed by Mohri and Hashimoto [21]. As the figure shows, the boosted score allows for higher claim retention across all datasets (mean retention of 39% vs. 24% for the boosted vs. unboosted scores).
Our second method, which we call level-adaptive conformal prediction, allows the validity of the conformal output to depend on characteristics of the queried prompt. In our LLM experiments, we adapt the level, i.e., the claimed probability of correctness, individually to each prompt in order to ensure that issued outputs retain at least $70\%$ of the original set of sub-claims. For example, in Figure 1, we prompt GPT-3.5-Turbo to output a response to a question from the MedicationQA dataset [6]. Outputting a filtered response that achieves the stated factuality criterion with probability $90\%$ requires near-complete censorship, but by relaxing the level to $63\%$ using our method, we can preserve almost the entire response.
@@
Figure 2 displays the advantages of our approach to this problem. First, the left panel of Figure 2 verifies that the level-adaptive probabilities we report are empirically well-calibrated. Second, the right panel of Figure 2 quantitatively demonstrates the improved claim retention of our method and verifies that for each dataset included in the MedLFQA benchmark [13] our level-adaptive conformal prediction retains at least 70% of the original output’s claims in most examples. Finally, by combining our level-adaptive and conditional boosting methods, we retain most claims and output non-trivial guarantees of response factuality; the left panel shows that the issued probabilities vary between $50$ and $85\%$. By contrast, while the fixed level method guarantees a $90\%$ probability of correctness, the method retains very little of the original LLM output.
@@
More concretely, (1) establishes that the issued probabilities are well-calibrated in the following sense: among similar prompts, the outputs that we claim to be factually correct with probability, say, between 70 and 80% will be actually factual between 70 and 80% of the time. In Section 3.1, we show how our framework can be adapted to guarantee that the LLM’s response satisfies other alignment targets beyond factual accuracy.
@@
We release a filtered version of the MedLFQA benchmark that removes some non-health-related prompts, the generated and parsed text used to run our experiments, as well as the notebooks used to produce the figures in this paper at github.com/jjcherian/conformal-safety. We also update our Python package for conditional conformal inference to support level-adaptive conformal prediction. This package is available to download at github.com/jjcherian/conditional-conformal and can be installed from PyPI.
@@ 2.1 Conformal prediction with conditional guarantees
To make this guarantee concrete, consider the case where :=$\mathcal{F}:=\{(\mathbbm{1}\{X\in G\})^{\top}\beta\mid\beta\in\mathbb{R}^{|\mathcal{G}|}\}$ for some set of subgroups $\mathcal{G}$. Then, (3) is exactly equivalent to group-conditional coverage, i.e., $\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1})\mid X_{n+1}\in G)=1-\alpha$ for all $G\in\mathcal{G}$.
To understand their construction, let $S(X,Y)\in\mathbb{R}$ denote a conformity score that measures how well $Y$ “conforms” to an estimate of its value. A typical choice in the regression setting is $S(X,Y)=|Y-\hat{\mu}(X)|$ for some fixed regression function $\hat{\mu}(\cdot)$; in the classification setting, we might choose $S(X,Y)=1/\hat{\pi}(Y\mid X)$ for some estimated conditional probabilities $\hat{\pi}$.
Gibbs et al. [10] estimate a high-probability upper bound for these scores by fitting an augmented quantile regression in which the unknown test score, $S(X_{n+1},Y_{n+1})$ is replaced by an imputed value $S$. Formally, let $\ell_{\alpha}(r):=(1-\alpha)[r]_{+}+\alpha[r]_{-}$ denote the pinball loss. Then, they define
@@
Since $\hat{C}(X_{n+1})$ can be mildly conservative, Gibbs et al. [10] also define a smaller randomized prediction set, $\hat{C}_{\text{rand.}}(X_{n+1})$ that achieves exact coverage. We will typically prefer to work with this set, but defer a detailed definition of its construction to the Appendix. As the following theorem shows, this randomized set achieves the guarantee stated in (3).
@@ 2.2 Conformal factuality
As discussed in the introduction, prediction sets are not suitable for many LLM use-cases. As an alternative, Quach et al. [25] and Mohri and Hashimoto [21] use conformal inference to develop filtering methods that remove false claims from an LLM’s output. We will focus specifically here on the work of Mohri and Hashimoto [21] since this is most similar to the new methods that we will propose in this article. Recall that $\mathbf{C}_{i}=\{C_{ij}\}_{j=1}^{k_{i}}$ denotes the claims made in an LLM’s response and $\mathbf{W}_{i}=\{W_{ij}\}_{j=1}^{k_{i}}$ denotes binary variables for which $W_{ij}=1$ indicates that the claim is true.
@@
where $\hat{\tau}$ is the $\frac{\lceil(1-\alpha)(n+1)\rceil}{n+1}$-quantile of the conformity scores,
@@
Mirroring the proof of split conformal prediction [23], Theorem 1 of [21] shows that if $\{({P}_{i},R_{i},\mathbf{C}_{i},\mathbf{W}_{i})\}_{i=1}^{n+1}$ are exchangeable, this method satisfies (5).
@@ 3.1 Generalization to alternative targets
Our first contribution in this paper will be to generalize the conditional framework of Gibbs et al. [10] to accommodate generic LLM alignment tasks. To do so, we extend the conformal risk control framework [3, 15] to provide high-probability control of a monotone loss with conditional guarantees.
More concretely, suppose that we are given a loss function $L(\hat{F}(\mathbf{C}_{i}),\mathbf{W}_{i})$ that measures the quality of the filtered output $\hat{F}(\mathbf{C}_{i})$ relative to the ground truth $\mathbf{W}_{i}$. Our goal will be to ensure that $\mathbb{P}(L(\hat{F}(\mathbf{C}_{n+1}),\mathbf{W}_{n+1})\leq\lambda)\geq 1-\alpha$, for some user-specified tolerance $\lambda>0$. For example, in the prior section $L(\cdot,\cdot)$ was the binary indicator that $\hat{F}(\mathbf{C}_{n+1})$ contains a false claim. More generally, $L(\cdot,\cdot)$ could measure the presence of toxic or discriminatory content in the response.
To incorporate general losses into the conditional conformal framework of Gibbs et al. [10], we require two assumptions. First, we assume that the method is always permitted to abstain from issuing a response and thus $L(\emptyset,\cdot)=0$. Second, we assume that the loss is monotone. Namely, for any sets of claims $\hat{F}_{1}(\mathbf{C}_{i})\subseteq\hat{F}_{2}(\mathbf{C}_{i})$, it must be the case that $L(\hat{F}_{1}(\mathbf{C}_{i}),W_{i})\leq L(\hat{F}_{2}(\mathbf{C}_{i}),W_{i})$.
With these assumptions, we extend the calibration procedure of Gibbs et al. [10] as follows. Let $X_{i}=X(P_{i},R_{i})$ denote a set of features computed using the prompt and response. Matching (6), we define $p(P_{i},C_{ij})$ to be a score measuring the quality of claim $C_{ij}$. We define the filtered set of claims by $\hat{F}(\mathbf{C}_{i})=F(\mathbf{C}_{i};\hat{\tau}):=\{C_{ij}:p(P_{i},C_{ij})\geq\hat{\tau}\}$, and the conformity score via
$S(\mathbf{C}_{i},\mathbf{W}_{i}):=\inf\{\tau\mid L(F(\mathbf{C}_{i};\tau),\mathbf{W}_{i})\leq\lambda\}.$
Our two assumptions (monotonicity of the loss and $L(\emptyset,\cdot)=0$) imply that $S(\mathbf{C}_{i},\mathbf{W}_{i})$ is well-defined and equal to the minimum loss-controlling threshold. Finally, we set
@@
Let =$\mathcal{F}=\{\Phi(X)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$ denote any finite dimensional linear class. Assume that $\{\mathbf{D}_{i}\}_{i=1}^{n+1}$ are exchangeable, that solutions to (7) and its dual are computed symmetrically in the input data, $L(\cdot,\cdot)$ is monotone in its first argument, and $L(\emptyset,\cdot)=0$. Then, for all $f\in\mathcal{F}$,
@@
As above, we can make this guarantee concrete by choosing $\mathcal{F}$ to be a linear combination of group indicators specifying the topic of the prompt, i.e., =$\mathcal{F}=\{(\mathbbm{1}\{X\in G\})^{\top}\beta\mid\beta\in\mathbb{R}^{|\mathcal{G}|}\}$. For any finite set of groups $\mathcal{G}$, this choice of $\mathcal{F}$ yields the high-probability guarantee,
@@
Like other conformal guarantees, the result of Theorem 3.1 only holds marginally over the calibration data, $\{\mathbf{D}_{i}\}_{i=1}^{n}$. Nevertheless, Gibbs et al. [10] observe that the calibration-conditional validity of their method concentrates around the nominal level as the calibration set size grows.
@@ 3.2 Level-adaptive conformal prediction
In addition to adapting the filtering cutoff to $X_{n+1}$, it may also be beneficial to adapt the desired error probability $\alpha$ at test time. [^3] Although our methods’ implementations differ, we remark here that the level-adaptive guarantee is inspired by a similar result in Zhang and Candès [33]. In this manuscript, we adjust $\alpha$ to ensure sufficient claim retention for each test output. In other scenarios, we might set a stricter $\alpha$ for high-stakes prompts and a more lenient one otherwise. For now, suppose that we are given some desired level function $\alpha(\cdot)$, and our goal is to ensure that $L(\hat{F}(\mathbf{C}_{n+1}),\mathbf{W}_{n+1})\leq\lambda$ with probability $1-\alpha(X_{n+1})$.
Recall that in the previous section, we use the pinball loss, $\ell_{\alpha}(\cdot)$ to learn the $1-\alpha$ quantile of $S(\mathbf{C}_{i},\mathbf{W}_{i})$. To control the error rate adaptively, here we will use a data-dependent loss for the $i$-th point, $\ell_{\alpha(X_{i})}(\cdot)$. Then, similar to the previous section, we define
@@
and filter at the cutoff $\hat{\tau}_{\text{l.a.}}(X_{n+1}):=\sup\{S:S\leq g_{S}(X_{n+1})\}$. As in the previous section, it is more convenient to work with a slightly smaller randomized cutoff, $\hat{\tau}_{\text{l.a., rand.}}$. The following theorem shows that our previous guarantee for fixed $\alpha$ extends to this new setting.
@@
While this choice of function class elicits an interpretable guarantee, it can be quite large in practice. This can slow down the computation of the conformal cutoff and reduce claim retention. Alternatively, this class might be chosen to exactly satisfy other popular (and more efficient) approximations to multicalibration such as low-degree multicalibration [11]. In Section B.2, we briefly explore the consequences of choosing an insufficiently complex function class in a synthetic regression example. A more substantial exposition of these trade-offs can be found in Gibbs et al. [10].